A. Write the missing term/s in the given series of numbers.

1. Find the next three terms in: 14, 10, 6, 2, ? , ? , ? .

2. Find the

th 35

term in the arithmetic sequence


3, 9, 15, 21, ? .


3. Find the next two terms in the sequence:


5, 2, 8, 3, 11, 4, 14, 5, 17, 6, ? , ? .


4. Find the next number:


2, 3, 4, 6, 6, 9, 8, ? .


5. What should become in place of x in the following series:


2, 2, 4, 12, 48, 240, x?


6. Identify the next number in the following series


2, 8, 26, 62, 122, 212, ? .


7. What is the next number in the pattern


4, 12, 36, 108, ? .


1
Expert's answer
2022-02-21T17:10:40-0500

A.

1. "-14, -10, -6, -2, \textcolor{blue}{2,4,8},"


2. "a_1=3, d=9-3=6"


"a_{35}=a_1+d(35-1)=3+6(35-1)=207"



3.

Consider two sequences


"5,8,11,14,17,20,23,..."

"2,3,4,5,6,7,8,..."

Hence we have the sequence


"5, 2,8, 3, 11, 4,14, 5,17,6, \textcolor{blue}{20,7},..."

4.

Consider two sequences


"2,4,6,8,10,..."

"3,6,9,12,..."

Hence we have the sequence


"2, 3,4,6, 6, 9, 8, \textcolor{blue}{12},..."


5.

We see


"2\cdot 1=2, 2\cdot 2=4, 4\cdot 3=12, 12\cdot 4=48,"

"48\cdot 5=240, 240\cdot 6=1440"

Hence "x=1440"


6.

"2\ \ \ \ \ \ \ 8\ \ \ \ \ \ \ 26\ \ \ \ \ \ \ 62\ \ \ \ \ \ \ 122\ \ \ \ \ \ \ 212\ \ \ \ \ \ \ \ 338\\"

"6\ \ \ \ \ \ \ \ 18\ \ \ \ \ \ \ \ 36\ \ \ \ \ \ \ \ 60 \ \ \ \ \ \ \ \ 90 \ \ \ \ \ \ \ \ 126\ \ \ \ \\"


"12\ \ \ \ \ \ \ \ 18 \ \ \ \ \ \ \ \ 24\ \ \ \ \ \ \ \ 30\ \ \ \ \ \ \ \ 36\ \ \ \ \ \\"

"6 \ \ \ \ \ \ \ \ \ \ 6\ \ \ \ \ \ \ \ \ \ 6\ \ \ \ \ \ \ \ \ \ 6\ \ \ \ \ \\ \\"

Hence we have the sequence


"2,8, 26, 62, 122, 212, \textcolor{blue}{338},..."



7.


"4\times 3=12, 12\times 3=36, 36\times 3=108,"

"108\times 3=324"

Hence we have the sequence


"4,12, 36, 108, \textcolor{blue}{324},..."